Three charges: Q1, Q2, and Q3 are located on a straight line. The charge Q3 = 6.75e-6 C is movable along this line. The charges Q1 = -8.75e-6 C and Q2 = 2.80e-6 C are fixed at their positions, with Q2 at the distance d = 0.436 m to the right of Q1. For what position of Q3 (with respect to Q1) is the net force on Q3 - due to Q1 and Q2 - equal zero? Convention: use the plus sign for Q3 to the right of Q1. (in m)
Answer #1

01 Q3 F23 0.436 m F12 (0.436+x)

F23 = F13

ke | Q2 | | Q3 | / r232 = ke | Q1 | | Q3 | / r132

| Q2 | | Q3 | / r232 = | Q1 | | Q3 | / r132

(2.8 x 10-6 C) (6.75 x 10-6 C) / (0.436 m)2 = (8.75 x 10-6 C) (6.75 x 10-6 C) / [(0.436 m) + x]2

(2.8) [(0.436 m) + x]2 = (8.75) (0.436 m)2

(2.8) [(0.190096 m2) + (0.872 m) x + x2] = (1.66334 m2)

(0.5322688 m2) + (2.4416 m) x + (2.8) x2 = (1.66334 m2)

(2.8) x2 + (2.4416 m) x - (1.1310712 m2) = 0

it's an quadratic equation and we get -

x = - 1.20675 m        &    x = 0.334746 m

Distance from Q1 = x13

we know that,   x13 = (0.436 m) + x

Firstly, we have

x13 = (0.436 m) + (-1.20675 m)

x13 = - 0.77075 m

Secondly, we have

x13 = (0.436 m) + (0.334746 m)

x13 = 0.770746 m